# Towards Fast Inference: Exploring and Improving Blockwise Parallel Drafts

### v1

@@ B.2 Future Work
  • The intrinsic compatibility of our lattice rescoring method with alternative sampling strategies presents a fertile ground for future inquiries. This delineation not only enriches the discourse within efficient language model inference but also sets a trajectory for subsequent empirical endeavors.
  • Scaling the BPD model for compatibility with larger-scale LLMs, ensuring it remains effective and efficient as language models evolve.
@@
  • Using the sequential entropy ordering of heads (Figure 3) as a possible halting condition during BPD head training, or to inform how a rescoring LM should be interpolated with the BPD lattice weights.
@@ Appendix C Experiment Details
C.1 Training Objective for BPD LMs
We minimized the following loss function to train BPD LMs:
@@ C.1 Training Objective for BPD LMs
where $x_{1\ldots i}$ is the token sequence up to position $i$, $y_{i+h}$ is the ground truth token at position $i+h$, and $p(y_{i+h}|x_{1\ldots i})$ is the probability of observing token $y_{i+h}$ given the sequence $x_{1\ldots i}$ under the BPD LM. We trained all models in this work with $\lambda_{h}=1$. We leave tuning these hyperparameters, improving the block efficiency and quality of the BPD LM, as future work.
@@ C.2 Neural Model Details
Each neural rescoring LM is a decoder-only transformer with learned absolute positional embeddings and twelve self-attention heads at each layer. The key architecture hyperparameters are given in Table 6. Aside from scale, the only difference between the BPD model and neural rescoring models is the addition of the feedforward neural networks and eight additional block prediction heads. Note that the number of parameters for each of these models also includes the embedding table.
Each model was pretrained on the English C4 corpus for 200K iterations with a batch size of $2^{20}\approx 1M$ tokens per batch. Dropout was not applied. For the BPD model, all heads were trained jointly.
@@
All n-gram  LMs in this work are Katz backoff n-gram LMs (Katz, 1987) fit on the train split of the GPT3 subword-tokenized English C4 corpus with n-gram order $\in\{2,4\}$. We apply entropy pruning (Stolcke, 2000) to reduce model size to a maximum of 100 million n-grams per model, and ensure that each trigram is observed at least twice and each 4-gram is observed at least four times. Preprocessing of the text is identical to that used to train neural LMs.
@@ C.4 Datasets
  • LAMBADA (LAnguage Modeling Broadened to Account for Discourse Aspects): A collection of narrative passages designed to test the understanding of long-range dependencies in language models, where the task involves predicting the last word of a passage based on the full context (Paperno et al., 2016).
  • SQuAD V1 (Stanford Question Answering Dataset): A reading comprehension dataset that features questions based on Wikipedia articles, with answers located within the text (Rajpurkar et al., 2016).
  • CNN/DailyMail: This dataset includes news articles paired with human-written summaries, mainly used to evaluate the summarization capabilities of language models, particularly in abstractive summarization (Hermann et al., 2015).
  • SAMSum (Semi-Automatic Machine Summarization): Focuses on abstractive summarization using news articles and machine-generated summaries, testing models’ abilities to refine and improve existing summaries (Gliwa et al., 2019).
  • MultiNews: Comprises news articles from diverse sources for abstractive summarization tasks, evaluating models on handling different writing styles and topics (Fabbri et al., 2019).
  • XSUM: Contains scientific documents and summaries, challenging language models to process complex scientific information and language (Narayan et al., 2018).
  • NewsRoom: A dataset of news articles aimed at assessing the factual accuracy and information extraction capabilities of models in generating summaries (Grusky et al., 2018).
All datasets were tokenized using the 50,257 GPT3 subword vocabulary (Brown et al., 2020).
@@
  • SQuAD: ”question: [question] context: [context]”
  • CNN/DailyMail: ”summarize: [text]”
@@
  • MultiNews: ”Write a summary based on this article: [text]”
  • XSUM: ”Summarize: [text]”
  • NewsRoom: ”Please write a short summary for the following article: [title] [text]”
@@ Appendix E Interpolation Weights Tuned per Task
Note that for tasks where lattice rescoring was unhelpful, the interpolation weight, $\alpha$ is tuned to place much higher weight on the BPD logits (Table 8). This is a signal that the rescorer does not provide additional information over the original BPD heads.
@@ Appendix F Local Rescoring Impact on Block Efficiency
Task Dataset Base 2-gram 3-gram 4-gram neural-32M neural-61M neural-94M Oracle (k=2) Oracle (k=16)
@@
Table 9 reveals the impact of different rescoring methods on the block efficiency of the BPD lattice, offering valuable insights into their effectiveness across diverse tasks and models, supporting the investigations in section 7.
  • Limited improvement for high baselines: For tasks with already high initial block efficiency (LAMBADA, CNN/DailyMail), rescoring offers minimal or even negative changes in block efficiency compared to the baseline BPD system. This suggests that for well-calibrated models with hierarchical confidence ordering, the standard BPD already achieves significant speed improvements, leaving limited room for further gains through rescoring.
  • Efficacy for poor baselines: In tasks with less calibrated predictions (SQuAD V1, XSUM, NewsRoom), rescoring using both N-gram and neural language models demonstrably improves block efficiency. Notably, neural rescoring with larger models (61M and 94M parameters) achieves the highest efficiency gains in these tasks, reaching up to 19.44% improvement in NewsRoom. These results highlight the potential of rescoring to refine predictions and enhance efficiency for models exhibiting calibration issues.
  • Task-specific effectiveness: The level of improvement from rescoring varies across different summarization tasks (MultiNews, XSUM, NewsRoom). While all show positive responses, NewsRoom exhibits the most significant gains, suggesting that the effectiveness of rescoring can be task-dependent.
@@
Appendix G Local Rescoring via Neural Models vs Parallel BPD via Global n-gram Rescoring
This section supports the detailed explanations on two algorithms in section 7: local rescoring via neural models and parallel BPD via global n-gram rescoring.
  • Local rescoring via neural models typically refers to a process where predictions are adjusted at each individual step of the sequence generation. A local rescoring algorithm takes the initial predictions from the model and refine them using additional information or a secondary model.
  • Parallel BPD via global n-gram rescoring, on the other hand, implies a more comprehensive approach where the entire sequence or multiple sequences are considered together. The ‘global’ aspect means that the rescoring process takes into account the whole sequence’s context to find the optimal path, rather than making decisions based on individual tokens. Parallel batched processing indicates that multiple sequences are being evaluated and rescored in one go, which is often more efficient. n-gram rescoring uses the statistical properties of n-grams to predict the probability of a sequence, allowing for the consideration of the sequence as a whole rather than as isolated predictions.
Appendix H Ablation on the Number of Heads in the BPD Model
@@
Appendix I Efficiency of BPD Rescoring
To enhance our understanding of BPD rescoring within the realm of contemporary deep learning hardware environments, we present an in-depth examination focused on TPU/GPU utilization and the overhead incurred by n-gram  rescoring. This analysis is divided into two parts: (1) an analysis of BPD rescoring through the lens of TPU/GPU utilization, and (2) empirical benchmarks of n-gram  lattice rescoring. The major takeaways are as follows.
@@ Appendix I Efficiency of BPD Rescoring
A critical factor in the performance of deep learning applications is the efficient management of memory bandwidth between High Bandwidth Memory (HBM) and Static Random Access Memory (SRAM) (Dao et al., 2022). Increasing the block efficiency via BPD lattice rescoring reduces the average per token parameter and key-value cache I/O that needs to be communicated from HBM to SRAM.
@@ I.1 TPU/GPU Utilization
Table 11: Comparative analysis of per decoded token efficiency metrics across BPD rescoring methods and the traditional autoregressive LM (batch size=1). This table showcases the average block efficiency, parameter I/O, key-value (KV) cache I/O at varying sequence lengths, and FLOPS—evaluated on a per-token basis under the condition of batch size 1.
Component Autoregressive Base BPD 4-gram BPD Neural-61M BPD Parallel Base BPD Parallel 4-gram BPD
@@
The BPD rescoring variants achieve significant reductions in Parameter I/O and KV Cache I/O compared to autoregressive decoding, suggesting BPD methods’ ability to reducing inference times by mitigating the primary latency bottleneck—memory bandwidth. Advances in TPU/GPU architecture ensure that an increase in FLOPS per token has a minimal effect on latency, confirming our strategy’s capacity to navigate the complexities of memory bandwidth efficiently.
@@
A consistent decrease in memory bandwidth utilization across BPD LMs, including those leveraging LM rescoring and parallel processing strategies, illustrates our approach’s contribution to accelerating inference speed. This underscores the practicality and applicability of our enhancements in promoting more efficient language model inference within state-of-the-art computational frameworks.
I.2 Overhead of n-gram Rescoring
While the majority of computational efforts in BPD rescoring are dedicated to TPU/GPU utilization, the implementation of n-gram rescoring introduces additional overheads. These are primarily attributed to CPU computations and the data transfer between the CPU and HBM. This section provides a comprehensive examination of these overheads, drawing on benchmarks from rescoring experiments with a 4-gram C4 LM.
@@ I.2 Overhead of n-gram Rescoring
We conducted benchmarks on rescoring lattices with a 4-gram C4 LM of $\approx$100M n-grams. The average latency observed across 10 runs for different numbers of the shortest paths is summarized in the Table 12:
@@
Notably, rescoring with a large 4-gram LM averages less than 2 milliseconds for extracting up to 16 globally-best paths, despite the lattice containing approximately 4.29 billion possible paths. In our initial experiments, increasing the size of the n-gram  LM had little effect on n-best rescoring latency, indicating that enhancements leading to higher block efficiency will incur little additional latency, provided the LM fits within DRAM.
Latency is predominantly influenced by lattice size, particularly the number of top-k tokens per time step and the number of time steps, as depicted in Table 13:

### v2

@@ B.2 Future Work
  • The intrinsic compatibility of our lattice rescoring method with alternative sampling strategies presents fertile ground for future work. This delineation not only enriches the discourse within efficient language model inference but also sets a trajectory for subsequent empirical endeavors.
  • Scaling the blockwise parallel LM for compatibility with larger-scale LLMs, ensuring it remains effective and efficient as language models evolve.
@@
  • Using the sequential entropy ordering of heads (3(b)) as a possible halting condition during block draft head training, or to inform how a rescoring LM should be interpolated with the block lattice weights.
@@ Appendix C Experiment Details
C.1 Training objective for blockwise parallel LMs
We minimized the following loss function to train blockwise parallel LMs:
@@ C.1 Training objective for blockwise parallel LMs
where $x_{1\ldots i}$ is the token sequence up to position $i$, $y_{i+h}$ is the ground truth token at position $i+h$, and $p(y_{i+h}|x_{1\ldots i})$ is the probability of observing token $y_{i+h}$ given the sequence $x_{1\ldots i}$ under the blockwise parallel LM. We trained all models in this work with $\lambda_{h}=1$. We leave tuning these hyperparameters, improving the block efficiency and quality of the blockwise parallel LM, as future work.
@@ C.2 Neural Model Details
Each neural rescoring LM is a decoder-only transformer with learned absolute positional embeddings and twelve self-attention heads at each layer. The key architecture hyperparameters are given in Table 6. Aside from scale, the only difference between the blockwise parallel LM and neural rescoring models is the addition of the feedforward neural networks and eight additional block prediction heads. Note that the number of parameters for each of these models also includes the embedding table.
Each model was pretrained on the English C4 corpus for 200K iterations with a batch size of $2^{20}\approx 1M$ tokens per batch. Dropout was not applied. For the blockwise parallel LM, all heads were trained jointly. The pretraining for the blockwise parallel LMs took about 47 hours on 128 TPUv3 units.
@@
All n-gram  LMs in this work are Katz backoff n-gram LMs [19] fit on the train split of the GPT3 subword-tokenized English C4 corpus with n-gram order $\in\{2,4\}$. We apply entropy pruning [39] to reduce model size to a maximum of 100 million n-grams per model, and ensure that each trigram is observed at least twice and each 4-gram is observed at least four times. Preprocessing of the text is identical to that used to train neural LMs.
@@ C.4 Datasets
  • LAMBADA (LAnguage Modeling Broadened to Account for Discourse Aspects): A collection of narrative passages designed to test the understanding of long-range dependencies in language models, where the task involves predicting the last word of a passage based on the full context [29].
  • SQuAD V1 (Stanford Question Answering Dataset): A reading comprehension dataset that features questions based on Wikipedia articles, with answers located within the text [34].
  • CNN/DailyMail: This dataset includes news articles paired with human-written summaries, mainly used to evaluate the summarization capabilities of language models, particularly in abstractive summarization [15].
  • SAMSum (Semi-Automatic Machine Summarization): Focuses on abstractive summarization using news articles and machine-generated summaries, testing models’ abilities to refine and improve existing summaries [12].
  • MultiNews: Comprises news articles from diverse sources for abstractive summarization tasks, evaluating models on handling different writing styles and topics [11].
  • XSUM: Contains scientific documents and summaries, challenging language models to process complex scientific information and language [28].
  • NewsRoom: A dataset of news articles aimed at assessing the factual accuracy and information extraction capabilities of models in generating summaries [13].
All datasets were tokenized using the 50,257 GPT3 subword vocabulary [3].
@@
  • SQuAD: "question: [question] context: [context]"
  • CNN/DailyMail: "summarize: [text]"
@@
  • MultiNews: "Write a summary based on this article: [text]"
  • XSUM: "Summarize: [text]"
  • NewsRoom: "Please write a short summary for the following article: [title] [text]"
@@ Appendix E Interpolation weights tuned per task
Note that for tasks where lattice rescoring was unhelpful, the interpolation weight, $\alpha$ is tuned to place much higher weight on the block draft logits (Table 8). This is a signal that the rescorer does not provide additional information over the original block draft heads.
@@ Appendix F Local rescoring impact on block efficiency
Task Dataset Baseline Global rescoring Local rescoring Oracle (k=2) Oracle (k=16)
BPD 2-gram BPD 3-gram BPD 4-gram BPD neural-32M BPD neural-61M BPD neural-94M BPD
@@
Table 9 reveals the impact of different rescoring methods on the block efficiency of the block lattice, offering valuable insights into their effectiveness across diverse tasks and models, supporting the investigations in subsection 7.1.
  • Limited improvement for high baselines: For tasks with already high initial block efficiency (LAMBADA, CNN/DailyMail), rescoring offers minimal or even negative changes in block efficiency compared to the baseline BPD system. This suggests that for tasks where standard BPD already achieves significant speed improvements, there is limited room for further gains through rescoring.
  • Efficacy for poor baselines: In tasks with lower initial block efficiency (SQuAD V1, XSUM, NewsRoom), rescoring using both n-gram and neural language models results in increased block efficiency. Notably, neural rescoring with larger models (61M and 94M parameters) achieves the highest efficiency gains in these tasks, reaching up to 19.44% improvement in NewsRoom. These results highlight the potential of rescoring to refine predictions and enhance efficiency for models exhibiting calibration issues.
  • Task-specific effectiveness: The level of improvement from rescoring varies across different summarization tasks (MultiNews, XSUM, NewsRoom). While all show positive responses, NewsRoom exhibits the largest gains, suggesting that the effectiveness of rescoring is task-dependent.
@@
Appendix G Ablation on the number of heads in the blockwise parallel LM
@@
Appendix H Practical efficiency of rescoring block drafts
To enhance our understanding of block rescoring within the realm of contemporary deep learning hardware environments, we present an in-depth examination focused on TPU/GPU utilization and the overhead incurred by n-gram  rescoring. This analysis is divided into two parts: (1) an analysis of block rescoring through the lens of TPU/GPU utilization, and (2) empirical benchmarks of n-gram  lattice rescoring. The major takeaways are as follows.
@@ Appendix H Practical efficiency of rescoring block drafts
A critical factor in the performance of deep learning applications is the efficient management of memory bandwidth between High Bandwidth Memory (HBM) and Static Random Access Memory (SRAM) [8]. Increasing the block efficiency via the block lattice rescoring reduces the average per token parameter and key-value cache I/O that needs to be communicated from HBM to SRAM.
@@ H.1 Hardware utilization
Table 11: Comparative analysis of per decoded token efficiency metrics across block rescoring methods and the standard autoregressive LM (batch size=1). This table shows the average block efficiency, parameter I/O, key-value (KV) cache I/O at varying sequence lengths, and FLOPS—evaluated on a per-token basis with batch size 1.
Component Autoregressive Base BPD 4-gram BPD Neural-61M BPD 16-best 0-gram BPD 16-best 4-gram BPD
@@
The block rescoring variants achieve significant reductions in Parameter I/O and KV Cache I/O compared to autoregressive decoding, suggesting BPD methods’ ability to reducing inference times by mitigating the primary latency bottleneck—memory bandwidth. Advances in TPU/GPU architecture ensure that an increase in FLOPS per token has a minimal effect on latency, confirming our strategy’s capacity to navigate the complexities of memory bandwidth efficiently.
@@
A consistent decrease in memory bandwidth utilization across blockwise parallel LMs, including those leveraging LM rescoring and parallel processing strategies, illustrates our approach’s contribution to accelerating inference speed. This underscores the practicality and applicability of our enhancements in promoting more efficient language model inference within state-of-the-art computational frameworks.
H.2 Overhead of n-gram rescoring
While the majority of computational efforts in block rescoring are dedicated to TPU/GPU utilization, the implementation of n-gram rescoring introduces additional overheads. These are primarily attributed to CPU computations and the data transfer between the CPU and HBM. This section provides a comprehensive examination of these overheads, drawing on benchmarks from rescoring experiments with a 4-gram C4 LM.
@@ H.2 Overhead of n-gram rescoring
We conducted benchmarks on rescoring lattices with a 4-gram C4 LM of $\approx$100M n-grams. The average latency observed across 10 runs for different numbers of the shortest paths is summarized in Table 12.
@@
Notably, rescoring with a large 4-gram LM averages less than 2 milliseconds for extracting up to 16 globally-best paths, despite the lattice containing approximately 4.29 billion possible paths. In our initial experiments, increasing the size of the n-gram  LM had little effect on n-best rescoring latency, indicating that improvements to rescoring LM quality will incur little additional latency, provided that the rescoring LM fits within DRAM.
Latency is predominantly influenced by lattice size, particularly the number of top-k tokens per time step and the number of time steps, as depicted in Table 13.